OpenStudy (anonymous):

Rachel deposited $5,960.32 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

i'll give you the formula - try and calculate this value of t (time in years) yourself: 9000 = 5960.32 ( 1 + 0.042/2) ^ 2t 9000 = 5960.32* (1.021)^2t 1.021^2t = 9000 / 5960.32 2t * ln 1.021 = ln (9000/5960.32) t = [ ln (9000/5960.32) / ln 1.021 ] / 2

OpenStudy (anonymous):

i dont have a calculator :p

OpenStudy (anonymous):

B = D ( 1 + r/n ) ^(nt) 9,000 = 5960.32 ( 1 + 4.2%/2) ^(2t) 1 + .042/2 = 1 + .021 = 1.021 9000/ 5960.32 = 1.51 -> 1.021^(2t) = 1.51 -> 2t ln 1.021 = ln 1.51 => t = ln 1.51/ 2 ln 1.021 = 10 yrs

OpenStudy (anonymous):

yes thats what i got 10 years

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Thanks you for your verification :)

OpenStudy (anonymous):

yw